OpenStudy (anonymous):

Factor: x^2 - 144 a. (x-12)^2 b. (x+12)(x-12) c. (x+2)(x-72) d. (x-16)(x+9)

OpenStudy (anonymous):

\[ a^2 - b^2= (a-b)(a+b) \]

OpenStudy (anonymous):

\[ x^2-144= x^2 -12^2 \]

OpenStudy (anonymous):

\[ x^2-144= x^2 -12^2=(x-12)(x+12) \]

OpenStudy (anonymous):

Thank you